d only on number 9 and b in number 7
(1096) Problem 9: Suppose you have a 9 00 V battery, a 2.4 SaF capacitor, and a 2.05 yF capacnor Completod - 7Partial Completed 0-94.05 10 10h, Completta l- 2596 Part (d) Find Grade Smary
Click here for 50% Part (a) what s the total charge on sphere two Q2 after they are connected, in 2:-1616 X Incerrect Feedback is avaulable Grade Summary Sabmiios 7 Partial 8 Completed cotanO asinacos0 10 Completed Hint hint. Hints
Answer #1

1)

Energy, E = 0.5 CV^2

E = 0.5 x 1.85 x 10^-6 x 9^2 = 7.49 x 10^-5 J

2)

Density, s = 1.616/(4 pi x 0.38^2) = 0.89 x 10^-6 C/m^2

Comment in case any doubt please rate my answer .. ..